Differentiate the Inverse Trigonometric functions:

1. Find 𝑓′′(πœ‹) if 𝑓(π‘₯) = π‘Žπ‘Ÿπ‘π‘ π‘–π‘›(π‘π‘œπ‘ π‘₯).

2. If π‘π‘œπ‘‘^βˆ’1(π‘₯𝑦) βˆ’ π‘‘π‘Žπ‘›^βˆ’1(𝑦/π‘₯) = 0, find 𝑑𝑦/𝑑π‘₯.

3. β„Ž(𝑑) = π‘Žπ‘Ÿπ‘π‘π‘œπ‘  (π‘‘βˆ’1/𝑑+1).


1
Expert's answer
2020-12-21T17:31:03-0500

1) Here "f(x)=arcsin(cosx)"

Therefore "f'(x)=\frac{1}{\sqrt{1-cos^2x}}.\frac{d}{dx}(cosx)"

"=\frac{1}{\sqrt {sin^2x}}.(-sinx)"

"=\frac{(-sinx)}{sinx}"

"=-1"

"\therefore f''(x)=0"

Hence, "f''(\pi)=0"


2)Given that , "Cot^{-1}(xy)-tan^{-1}(\frac {y}{x})=0" ........."(1)"

Now differentiating both side of (1) with respect to "x" ,we get

"\implies \frac{(-1)}{1+x^2y^2}.\frac{d}{dx}(xy)-\frac{1}{1+(\frac{y}{x})^2}.\frac{d}{dx}(\frac{y}{x})=0"

"\implies" "\frac{1}{1+x^2y^2}.(y+x.\frac{dy}{dx})+(\frac{x^2}{x^2+y^2}).(\frac{x\frac{dy}{dx}-y}{x^2})=0"

"\implies" "\frac{y}{1+x^2y^2}+\frac{x\frac{dy}{dx}}{1+x^2y^2}+\frac{x\frac{dy}{dx}}{x^2+y^2}-\frac{y}{x^2+y^2}=0"

"\implies y.(\frac{1}{1+x^2y^2}-\frac{1}{x^2+y^2})+x\frac{dy}{dx}.(\frac{1}{1+x^2y^2}+\frac{1}{x^2+y^2})=0"

"\implies y.[\frac{x^2+y^2-x^2y^2-1}{(1+x^2y^2)(x^2+y^2)}]+x\frac{dy}{dx}.[\frac{x^2+y^2+x^2y^2+1}{(1+x^2y^2)(x^2+y^2)}]=0"

"\implies \frac {dy}{dx}=-(\frac {y}{x}).[\frac {x^2+y^2-x^2y^2-1}{x^2+y^2+x^2y^2+1}]"


3) Given ,"h(t)=arccos(\frac {t-1}{t+1})"

Then "h'(t)=\frac{-1}{\sqrt{1-(\frac {t-1}{t+1})^2}}.\frac{d}{dt}(\frac{t-1}{t+1})"

"=\frac{-1}{\sqrt{1-(\frac {t-1}{t+1})^2}}.[\frac {(t+1). \frac{d}{dt}(t-1)-(t-1).\frac{d}{dt}(t+1)}{(t+1)^2}]"

"=\frac {-(t+1)}{\sqrt{(t+1)^2-(t-1)^2}}. [\frac {(t+1). 1-(t-1).1}{(t+1)^2}]"

"=\frac {-(t+1)}{\sqrt{4t}}. [\frac {2}{(t+1)^2}]"

"=\frac{-(t+1)}{2\sqrt{t}}.[\frac{2}{(t+1)^2}]"

"=\frac{-1}{\sqrt{t}.(t+1)}"